# A scalable quantum-neural hybrid variational algorithm for ground state estimation

### v1

@@ Appendix G Notes on training of the parameters
The VQNHE and U-VQNHE system consist of two sets of parameters: variational parameters for the quantum circuits and the parameters for the classical neural network. All simulations in the research has been conducted by training the VQE first, and then training the neural network on top of the already trained VQE. The reason for that is that, possible as it may seem, joint training of the parameters is impossible in the case of VQNHE, or in general, quantum algorithms with neural post-processing. When the VQE output is fixed, the neural network evaluation is deterministic, and its optimization is stable, whereas when the output of the quantum circuit is not fixed, it yields deviations on every measurement process. For joint training, this is a critical limitation, as the VQE yields unstable outputs and the neural network struggles to find a way to be trained such that it consistently provides downward gradient. Moreover, repetitive evaluation of the quantum circuit is very costly, making joint training very inefficient.

### v2

@@ Appendix G Notes on training of the parameters
The VQNHE and U-VQNHE system consist of two sets of parameters: variational parameters for the quantum circuits and the parameters for the classical neural network. All simulations in the research have been conducted by training the VQE first, and then training the neural network on top of the already trained VQE. The reason for that is that, possible as it may seem, joint training of the parameters is impossible in the case of VQNHE, or in general, quantum algorithms with neural post-processing. When the VQE output is fixed, the neural network evaluation is deterministic, and its optimization is stable, whereas when the output of the quantum circuit is not fixed, it yields deviations on every measurement process. For joint training, this is a critical limitation, as the VQE yields unstable outputs and the neural network struggles to find a way to be trained such that it consistently provides downward gradient. Moreover, repetitive evaluation of the quantum circuit is very costly, making joint training very inefficient.